# Oort Cloud Bombardment by Dark Matter Jeremy Mould

### v1

@@ 0 Introduction
Other signs of DM activity in the solar system have been suggested. De Rocco (2025) notes that DM in the form of macroscopic composites is largely unconstrained at masses of 1011 – 1017g. In this mass range, DM may collide with planetary bodies, depositing energy and leaving dramatic surface features that remain detectable on geological timescales. He suggests that Ganymede, the largest Jovian moon, provides a prime target to search for DM impacts due to its differentiated composition and Gyr-old surface. There are also searches for microscopic dark matter, such as WIMP capture in the Sun. Stothers (1984) was first to connect DM and comets, after the appearance of heavy halo models of the Galaxy (e.g. Caldwell & Ostriker 1981).
The collision rate for objects incident on the solar system is v n , where is the geometric cross section for the solar system, v is their velocity and n is their number density. The ratio of dark matter rate to stellar rate is nDM/n∗ vDM/v∗ which equals /mDM vDM divided by /m∗ v∗, where is density and m is mass. For mDM = 10-7 M⊙, the rate ratio is 107 0.01/0.1 220/30 for old disk stars that dominate the solar neighbourhood. So the DM bombardment rate is ten million times larger than the stellar rate. In fact, the conventional theory is not that stars perturb the Oort cloud. It is that the tidal field of the Galaxy perturbs the Oort cloud.
The nature of the DM remains one of the outstanding questions of 21st century astrophysics. Microlensing observations with the Rubin telescope could be prioritorized to discover subsolar mass objects (Romao, Croon & Godines 2025). Strong limits on fundamental particle DM are being set by physicists (Aalbers et al. 2023, Barberio et al. 2025). These may nevertheless be an important component of the DM, but they do not dislodge protocomets.
In the next section a toy model is developed to examine the effectiveness of lunar mass DM. In 2 results of these simulations are presented. They are discussed in 3. The conclusion is that within the mass range discussed macroscopic DM is a plausible contributor to the rate of comet production.
1 A toy model
The cumulative number of protocomets in the Oort cloud is given as N D-α, for D 2.8 km, with = 3.6 for larger D 2.8, and = 0.5 for smaller objects (Wajer et al. 2024). The Oort cloud contains 3.8 108 comets with D 10 km (Nesvorný 2018). Extending this to decimeter size (Vida et al. 2023) we obtain N = 1.5 1014, compared with 1012 – 1013 (Wiegert & Tremaine 1998). The Oort cloud is taken to be spherical, which is appropriate to its outer parts. The number of comets in the simulation is 250,000, compared with our adopted N = 4 1013 and so the fraction of 4 sr in the simulation is given by = 2.5 105/N, where is the half angle.
@@ 1 A toy model
The impact parameter, b, of the incoming object is a crucial parameter. I assume this lies with uniform probability within a sphere of radius b M AU, where M is the mass of the object in M⊙ units. For an impact velocity, v, this gives a momentum transfer, when the interaction with the protocomet is complete, of,
with m the mass of the perturbed particle, whose velocity is parallel to the vector from its original position to the impact parameter position. We then follow the orbits of 250,000 such particles with time intervals of 0.10 to 0.3 yr. The integration scheme was simply v = a dt, where a is the acceleration, followed by r = v dt, where v is the velocity. Simulations with different DM masses and other parameters are recorded in Table 1. Any protocomet reaching the inner solar system with radius 300 AU is noted. This choice of radius is on the large size, but the simulations are limited by small number statistics, and a markedly smaller radius would be prohibitive in computer time and not necessary for this first exploration of the hypothesis.
@@
There are two factors for which the simulation comet finding rate needs to be corrected, (1) the lost area factor in which DM particles falling outside the maximum adopted impact parameter are neglected (4 / N) and (2) the number of protocomets in the simulation, compared with N. With these corrections the predicted rate of entry of comets into the inner solar system222The Oort cloud is generally considered a spherical or almost spherical region, but it also has an inner, disc-shaped component that is more aligned with the ecliptic plane (the plane of the solar system). The outer Oort cloud is a vast, spherical halo extending far beyond the Kuiper Belt, while the inner Oort cloud is thought to be more of a donut shape and closer to the ecliptic. is
where n is the number recorded in the simulation in time t, and b7 = b 107. Figures 2 & 3 show the particle distribution and their velocity distribution shortly after the start of a simulation.
@@
1.1 Results
Figure 5 shows rate2 for a range of DM masses and maximum impact parameters, assuming that all the DM has mass M, i.e. fPBH = 1 in the case of PBH. The comet arrival rate in the inner solar system rate 10 per year for some maximum impact parameter for 10 M 10-5 M⊙. The rate would be brought down to order unity for fPBH = 0.1. The shallower Oort cloud distributions shown in Figure 1 would be expected to reduce the comet rate proportionally to the median radius. This was borne out in re-runs of # 19 and 20 with an r-3.5 density dependence.
2 Comet supply and demand
An equally fraught extrapolation to equation (3) in the present work is the estimation of the number of comets inside 300 AU from the known number inside 10 AU. The infall equation in the gravitational field is
for which, after integration, the steady state radial distribution in spherical symmetry is enclosed number n r3/2.
Figure 4 from the JPL small bodies database333https://ssd.jpl.nasa.gov shows n = 500 inside 10 AU; so n 82,000 inside 300 AU. With the rate found in the previous section, this can be furnished in a fraction of an Myr. But we need to consider the attrition rate.

### v2

@@ I Introduction
Other signs of DM activity in the solar system have been suggested. De Rocco (2025) notes that DM in the form of macroscopic composites is largely unconstrained at masses of 1011 – 1017g. In this mass range, DM may collide with planetary bodies, depositing energy and leaving dramatic surface features that remain detectable on geological timescales. He suggests that Ganymede, the largest Jovian moon, provides a prime target to search for DM impacts due to its differentiated composition and Gyr-old surface. There are also searches for microscopic dark matter, such as WIMP capture in the Sun. Stothers (1984) was first to connect DM and comets, after the appearance of heavy halo models of the Galaxy (e.g. Caldwell & Ostriker 1981). Brown, He & Unwin (2025) have explored the effect of PBH on exoplanet systems.
The collision rate for objects incident on the solar system is v n , where is the geometric cross section for the solar system, v is their velocity and n is their number density. The ratio of dark matter rate to stellar rate is (nDM/n (vDM/v∗) which equals (/m vDM divided by (/m v∗, where is density and m is mass. For mDM = 10-7 M⊙, the rate ratio is 107 0.01/0.1 220/30 for old disk stars that dominate the solar neighbourhood. So the DM bombardment rate is ten million times larger than the stellar rate. In fact, the conventional theory is not that stars perturb the Oort cloud. It is that the tidal field of the Galaxy perturbs the Oort cloud.
The nature of the DM remains one of the outstanding questions of 21st century astrophysics. Microlensing observations with the Rubin telescope could be prioritized to discover subsolar mass objects (Romao, Croon & Godines 2025). Strong limits on fundamental particle DM are being set by physicists (Aalbers et al. 2023, Barberio et al. 2025). These may nevertheless be an important component of the DM, but they do not dislodge protocomets.
In the next section a toy model is developed to examine the effectiveness of lunar mass DM. In II results of these simulations are presented. They are discussed in III. The conclusion is that within the mass range discussed macroscopic DM is a plausible contributor to the rate of comet production.
II A toy model
The cumulative number of protocomets in the Oort cloud is given as N D-α, for D 2.8 km, with = 3.6 for larger D 2.8 km, and = 0.5 for smaller objects (Wajer et al. 2024). The Oort cloud contains 3.8 108 comets with D 10 km (Nesvorný 2018). Extending this to decimeter size (Vida et al. 2023) we obtain N = 1.5 1014, compared with 1012 – 1013 (Wiegert & Tremaine 1998). The Oort cloud is taken to be spherical, which is appropriate to its outer parts. The number of comets in the simulation is 250,000, compared with our adopted N = 4 1013 and so the fraction of 4 sr in the simulation is given by = 2.5 105/N, where is the half angle.
@@ II A toy model
The impact parameter, b, of the incoming object is a crucial parameter. Unless a maximum value is adopted, beyond which the possibility of collision approaches zero, the computing time diverges, as all values of b are plausible out to the mean separation radius of protocomets. I therefore assume b lies with uniform probability within a sphere of radius b M AU, where M is the mass of the object in M⊙ units. For an impact velocity, v, this gives a momentum transfer, when the interaction with the protocomet is complete, of,
with m the mass of the perturbed particle, whose velocity is parallel to the vector from its original position to the impact parameter position. This is the impulse approximation, derived in full by Rickman (1976), the principal assumption being a high value of . We then follow the orbits of 250,000 such particles with time intervals of 0.10 to 0.3 yr. The appropriate time interval was monitored by noting when comets entering the inner solar system were double counted. The integration scheme was simply v = a dt, where a is the acceleration, followed by r = v dt, where v is the velocity. Simulations with different DM masses and other parameters are recorded in Table I. Any protocomet reaching the inner solar system with radius 300 AU is noted. This choice of radius is on the large size, but the simulations are limited by small number statistics, and a markedly smaller radius would be prohibitive in computer time and not necessary for this first exploration of the hypothesis.
@@
There are two factors for which the simulation comet finding rate needs to be corrected, (1) the lost area factor in which DM particles falling outside the maximum adopted impact parameter are neglected (4 /N) and (2) the number of protocomets in the simulation, compared with N. Each of these two multipliers contributes a factor of N to the numerator in equation (3). With these corrections the predicted rate of entry of comets into the inner solar system222The Oort cloud is generally considered a spherical or almost spherical region, but it also has an inner, disc-shaped component that is more aligned with the ecliptic plane (the plane of the solar system). The outer Oort cloud is a vast, spherical halo extending far beyond the Kuiper Belt, while the inner Oort cloud is thought to be more of a donut shape and closer to the ecliptic. is
where n is the number recorded in the simulation in time t, and b7 = b 107. Figures 2 & 3 show the particle distribution and their velocity distribution shortly after the start of a simulation. In Table I the number of comets observed in the inner solar system for each run is n. The maximum velocity attained by any of these is given in the centre column. It is clear that some comets, due to a decrease in orbital velocity, enter orbits of what are possibly observed as New comets. Depending on the direction of the primary black hole’s velocity vector, other comets may escape the solar system due to an increase in orbital velocity. Note values as high as 80 km s-1in Table 1. This is an important aspect of this problem, interesting in light of modern observations of interstellar comets.
@@
II.1 Results
Figure 4 shows rate2 for a range of DM masses and maximum impact parameters, assuming that all the DM has mass M, i.e. fPBH = 1 in the case of PBH. The comet arrival rate in the inner solar system rate 10 per year for some maximum impact parameter for 10 M 10-5 M⊙. The rate would be brought down to order unity for fPBH = 0.1. Microlensing constraints on PBH in this mass range are given in Table II from Green (2026), Sugiyama et al. (2026) and Key (2025). In the upper part of the mass range in Table 1 a mass fraction larger than 0.1 is ruled out by these measurements. The OGLE results are more stringent, but anomalous with respect to the other experiments. The ”asteroid window” (10-9 M⊙ and below) remains fully open because of diffraction limitations to microlensing.
| Survey | Duration | FoV | # stars | Detected | Mass range | fPBH |
| (years) | (deg2) | (million) | # events | M⊙ | ||
| MACHO | 5.7 | 14 | 12 | 13 | 3 10-7 to 5 10-4 | f 0.1 |
| EROS | 12 | 88* | 25* | 1 | 10-6 to 0.01 | f 0.1 |
| OGLE | 20 | 105† | 78.7† | 13 | 10-8 to 0.01 | f 0.01 |
| Subaru HSC | 39.3 hrs | 1.7 | 80 | 4 | 10-7 to 10-6 | 0.01 to 0.1 |
| AMDC | 5 nights | 3 | 4 | 1 | 10-8 to 10-5 | f 0.1 |
| *EROS-2, †OGLE-IV Data from Green (2026) FoV = field of view | ||||||
The shallower Oort cloud distributions shown in Figure 1 would be expected to reduce the comet rate proportionally to the median radius. This was borne out in re-runs of # 19 and 20 with an r-3.5 density dependence.
III Comet supply and demand
An equally fraught extrapolation to equation (3) in the present work is the estimation of the number of comets inside 300 AU from the known number inside 10 AU. Integrating d()/dr = –GM/r2, the infall equation in the gravitational field is
in the adopted units of velocity (30 km s-1) for which, after integration, the steady state radial distribution in spherical symmetry is enclosed number n r3/2.
Figure 5 from the JPL small bodies database333https://ssd.jpl.nasa.gov shows n = 500 inside 10 AU; so n 82,000 inside 300 AU. With the rate found in the previous section, this can be furnished in a fraction of an Myr. Further orbital parameters are in Figure 6. But we need to consider the attrition rate.